OpenStudy (anonymous):

Can someone give me examples on how you'd write functions for the following transformations using function notation using variables: translation of a units to the right and b units up reflection across the y-axis reflection across the x-axis rotation of 90 degrees counterclockwise about the origin, point O rotation of 180 degrees counterclockwise about the origin, point O rotation of 270 degrees counterclockwise about the origin, point O

OpenStudy (johnweldon1993):

Alright so say we have f(x) = x "translation of a units to the right and b units up " to move a graph up or down....you add or subtract a number 'b here' to the function... so f(x) = x + b To move it left or right you add or subtract a number 'a here' from the 'x' in the function...so f(x) = (x-a) + b

OpenStudy (johnweldon1993):

"reflection across the y-axis " for this...if we still have f(x) = x To reflect across the y-axis we need only put a '-' sign in front of the 'x' so it would be f(x) = -x

OpenStudy (johnweldon1993):

"reflection across the x-axis" For this f(x) = x We need to put the '-' sign in front of the function -f(x) = x and then simplify

OpenStudy (anonymous):

Thank you.

OpenStudy (johnweldon1993):

Still need the other ones too?

OpenStudy (johnweldon1993):

Well just in case I'll post them :) f(x) shifted upward a units = f(x) + a f(x) shifted downward a units = f(x) – a f(x) shifted left a units = f(x + a) f(x) shifted right a units = f(x – a) f(x) reflected about the x-axis = –f(x) f(x) reflected about the y-axis = f(-x) To rotate (A , B) 90 degrees counter clockwise -> (-B , A) To rotate (A , B) 180 degrees counter clockwise -> (-A , -B) To rotate (A , B) 270 degrees counter clockwise -> (B , -A)

OpenStudy (anonymous):

Yeah. Thank you so much!!!

OpenStudy (johnweldon1993):

No problem! :)

OpenStudy (anonymous):

For the rotations do you need to put "f(x) ="? o: